A UCLA researcher states that the life expectancy of mice can be extended by up to 45% when about 50% of the calories in their diet are reduced from the time they are weaned. The restricted diet is enriched to normal levels with vitamins and proteins. Assuming that it is known from previous studies that σ = 2.8 months, how many mice should be included in the sample to have 99% confidence that the expected half-life of the sample will be within 9.8 months from the population mean for all mice subject to the reduced diet?​​​​​​​
Answer #1

Given population standard deviation, \sigma = 2.8. Margin of error , e = 9.8

Confidence level = 99% = 100*(1-\alpha) %

or, \alpha = 1-99/100 = 0.01

or,1-\alpha/2 = 1-0.01/2 = 0.995

sample size, n = [Z(1-\alpha/2) * \sigma /e] 2

from Z-table, the closest value of Z corrensponding to probability of 0.995 is Z = 2.58.

** a table is given below for reference.

so, sample size = n = [2.58 * 2.8/9.8]2 = 0.54338

The sample size is 0.54338. Which is less than unity. We need to reduce the margin of error for a satisfactory sample size.

Z table -

了 .00 .01 .02 .03 104 .05 1,06 .07 108 009 0.0 diy .5000 .5040 0.1 1.5398 1.5438 0.2 1.5793 1.5832 0.3 .6179 .6217 0.4 1.6554